OpenStudy (anonymous):

Solve the given ODE. 1. (D^3+3D^2+3D+1)y=x^2 2.(D^4-2D^3-8D^2)y=6cosx 3.(D^3+1)y=9e^(x/2)

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

idk how to do dis sorry :(

OpenStudy (unklerhaukus):

1. \[ (D^3+3D^2+3D+1)y=x^2\]

OpenStudy (unklerhaukus):

to get the general solution, you need to find the homogenous solution, and the particular solution, first find the homogenous solution, to the homogenous equation: \[(D^3+3D^2+3D+1)y=0\]

OpenStudy (anonymous):

can you teach me how to get homogenous solution? :D

OpenStudy (unklerhaukus):

if we suppose \(y=e^{mx}\), \[D(y) = me^{mx}=my\]and\[D^2(y)=m^2e^{mx}=m^2y\] etc.

OpenStudy (unklerhaukus):

so \[(D^3+3D^2+3D+1)y=0\] will become \[(m^3+3m^2+3m+1)y=0\] and \(y=e^mx\) is not zero. so we can divide it out of the equation we get \[m^3+3m^2+3m+1=0\] can you solve this for values of \(m\)?

OpenStudy (unklerhaukus):

if if was a quadratic, we could use the quadratic formula, but alas, it is a cubic

OpenStudy (unklerhaukus):

can you factor it somehow?

OpenStudy (unklerhaukus):

\[(m-m_1)(m-m_2)(m-m_3)=0\]

OpenStudy (unklerhaukus):

(think of Pascal's triangle)

OpenStudy (anonymous):

is that the final answer?

OpenStudy (unklerhaukus):

nope, we have to find the solutions for \(m\) (all of them), then we will use them to get the homogenous solution to the differential equation.

OpenStudy (anonymous):

you mean solution set? how?

OpenStudy (unklerhaukus):

[ the particular solution is more difficult to find ]

OpenStudy (anonymous):

dont get mad. this is actually my boyfriend's homework. he is taking engineering course and i dont know how to solve this because im not into engineering. im so sorry. i just want to help him.

OpenStudy (anonymous):

how to find the particular solution?

OpenStudy (unklerhaukus):

probably best to use operator D, method

OpenStudy (unklerhaukus):

So far on our quest to find the homogenous solution, we have; \[m^3+3m^2+3m+1=0\] Which is a cubic equation. While cubics are usually horrid, this one is quite simple

OpenStudy (unklerhaukus):

Pascal's triangle looks something like this:  where terms below are made from the sum of the two number above

OpenStudy (unklerhaukus):

if we look at the \(\color{brown}{\textit{coefficients}}\) of the equation \[(\color{brown}1)m^3+(\color{brown}3)m2+(\color{brown}3)m+(\color{brown}1)=0\]

OpenStudy (anonymous):

what will i do with the coefficents? start factoring?

OpenStudy (unklerhaukus):

yes the \(m\) equation ends up factoring very nicely

OpenStudy (anonymous):

(3m-1) • (1) (m-3) • (m2)

OpenStudy (unklerhaukus):

A quadratic \[(x+1)^2=(x+1)(x+1) \\=x(x+1)+1(x+1)\\=x^2+2x+1 \\=(\color{brown}1)x^2+(\color{brown}2)x+(\color{brown}1)\] notice how the coefficients are a line in Pascal's triangle?